# Learning to Drive Anywhere with Model-Based Reannotation

### v1

@@ 4.2 Pre-Processing and Filtering
As shown on the leftmost side of Fig. 2, we use a classical state estimation pipeline to get better coarse robot pose estimates for FrodoBots-2k. We use a smoothing system based on a bidirectional Extended Kalman Filter (EKF) [40] to fuse raw actions with wheel speed measurements, GPS location, and compass heading (all of which are noisy) to get a smoothed estimate of the robot’s position. We also filter out data where the robot is paused for a long time to prioritize learning desirable behaviors. The cleaned and filtered data consists of approximately 700 hours of real-world navigation trajectories collected worldwide, which is still an order of magnitude larger than any currently available visual navigation dataset as shown in Table 1. While the EKF-based state estimation helps produce a less noisy action estimate [41], the signal remains too noisy for direct training. More details of the EKF design and its visualization are shown in the appendix.
4.3 Training Details
@@
We describe the training settings for both our short-horizon relabeling model, MBRA and long-horizon navigation policy, LogoNav.
@@
Short-horizon relabeling model: Since the robot system has $L$ steps system delay [42, 43] when operating remote robot via internet, we design our objective and network architecture to account for system delay to prevent overshooting or oscillating around target trajectories. Inspiring the previous works of model predictive control [44, 45], we consider the robotic states with the previous action commands $\{a_{i}\}_{i=-L\ldots-1}$ to genrate the actions $\{a_{i}\}_{i=0\ldots N-1}$.
@@
In training, we set the observation and action rate for trajectory sampling at 3 Hz for consistency with the GNM dataset. During training, we randomly select an image frame from the entire dataset as the current observation, and then randomly select a goal frame from up to $N_{g}=$ 20 steps (about 7 seconds) in the future. This short distance to the goal lets us learn precise labels to reannotate the action between $O_{c}$ and $O_{g}$. A detailed description of our MBL setup is available in the appendix.
@@
Long-horizon navigation policy: For long-horizon navigation, we use a larger $N_{g}=100$ to sample a goal position up to 33 seconds into the future. We reannotate actions with the short-horizon MBRA model to get high-quality action labels for the FrodoBots-2k dataset. This process yields action labels with a chunk size of $N=8$ steps. We train on the IL objective $J_{il}$ using the same parameters and settings as the short-horizon relabeling model otherwise.
Details of the network design and action and observation space definition are shown in the appendix.
5 Evaluation
@@
    Can we use MBRA to leverage action-free in-the-wild data?
@@ 5.1 Evaluation Setup
We concretely describe both short-horizon and long-horizon navigation tasks we evaluate our method on along with their associated baselines.
@@
Short-horizon navigation policy: Short-horizon navigation policy can navigate the robot toward a goal up to 3 meters away, so we use a topological memory to enable the robot to navigate to further goal positions, similar to other vision-based navigation approaches [10, 12, 9]. To collect this goal loop, we teleoperate the robot and record image observations at a fixed frame rate of 1 Hz. To deploy the policy, we start from the initial observation and continuously estimate the closest node as the current node at each time step, following [9, 27]. We feed the image from the next node as the goal image $O_{g}$ to our policy to compute the next action.
@@
Long-horizon navigation policy: Our long-horizon navigation policy can navigate to goals between 25-100 meters from the initial robot pose. We rely on GPS(outdoor) and tracking camera [46](inside) to get robot positions and specify goals. We evaluate longer trajectories by setting multiple subgoals at intervals of approximately 80 meters apart. At every time step, we calculate the current relative goal pose $p_{g}$ on the way to the next goal pose. When $|p_{g}|<5.0$ m, we consider the goal reached and update to the next subgoal for a longer trajectory.
We use the same robotic platform, Earth Rover Zero (ERZ), that was used to collect the FrodoBots-2k dataset for our main evaluation results and different robot platforms for cross embodiment analysis. See appendix for more details
5.2 Long-horizon Navigation Policy (LogoNav): GPS Goals
@@
Table 2: Evaluation of LogoNav on long-horizon pose-conditioned navigation tasks. “GS” and “COV” indicate the goal success rate and the coverage rate.
@@
NoMaD [27] GNM only - 0.333 0.471
Behavior Cloning EKF [41] 0.286 0.624
visual SLAM [26] 0.286 0.486
VPT [22] 0.095 0.314
multi-step VPT [22, 9] 0.619 0.757
@@
To answer Q1, we evaluate the long-horizon navigation policies trained with MBRA and five baselines: a NoMaD-like policy and behavior cloning (BC) policies trained on GNM + FrodoBots-2k with four different annotation approaches, 1) filtered action by EKF, 2) visual SLAM, 3) VPT [22], and 4) multi-step VPT [22, 9]. Details of the baselines are shown in the appendix. We select 7 outdoor locations and evaluate each policy 3 times for each goal. In Table 2, we show the goal success rate and the coverage rate for each method. The coverage rate is the ratio of the distance reached by the robot to the distance of the target goal pose before it fails. Our policy with MBRA shows stronger performance than the five baselines for both goal success rate and coverage rate. In Sec. 5.4, we conduct an investigation to analyze the advantageous gap of MBRA to answer Q3.
Figure 3 shows the third-person view at the start position and the robot trajectories on a bird-eye-view map in two scenes. Our policy distilled from MBRA actions was the only one to successfully navigate to the distant goal pose in both scenes, making a sharp left turn at the start to stay on path in case A. In contrast, both NoMaD and multi-step VPT could not execute this action, failing by colliding with bushes or requiring interventions to avoid falling down stairs. To show the capability of MBRA, we provide several subgoals, specified by latitude, longitude, and azimuth angle values, at intervals of approximately 80 meters, and evaluate LogoNav with MBRA on traversing these subgoals in two different scenes. As shown in Fig. 4, our navigation system with our policy enables us to navigate the robot toward a goal 300 meters away without collision, even in human-occupied spaces.
Moreover, we deployed LogoNav on two more robotic embodiments, including VizBot [47], a small Roomba-like robot, in an indoor setting, and the Unitree Go1 quadruped robot in an outdoor setting. We conducted 10 trials from up to 100 meters away in different challenging environments with some obstacles for each embodiment and method. We show the quantitative results in Table 3 and show the robotic behaviors in Fig. 5 and the supplemental videos. We achieve strong goal-reaching behavior with collision avoidance compared to the strongest baseline in Table 2, highlighting the policy’s generalization ability. Note that we apply the same policy in Table 2 and feed the generated actions without any adaptation. Action conversion is internally applied in each robot setup. Details of action space are mentioned in the appendix.
Table 3: Quantitative analysis with quadruped robot, Go1 and wheeled robot, VizBot.
@@
Behavior Cloning multi-step VPT 0.300 0.680 0.200 0.630
@@
Figure 5: Visualization of cross-embodiment analysis.
5.3 Training navigation policies on in-the-wild video with MBRA
Table 4: Evaluation of MBRA on action-free in-the-wild YouTube videos. “GS” and “SC” indicate the goal success rate and the subgoal coverage rate.
Dataset MBRA
@@ 5.3 Training navigation policies on in-the-wild video with MBRA
0.500 0.680
0.875 0.909
For Q2, we evaluate the capability of MBRA on different passive data sources, action-free in-the-wild video. We use the MBRA model to generate the action labels for the in-the-wild videos and train the short-horizon visual navigation policy conditioned on goal images, $\{a^{v}_{i}\}_{i=0\ldots N-1}=\pi^{v}(O_{c},O_{g})$. During training, we use the same objective $J_{il}$ to imitate the action labels generated by MBRA. We train two goal image-conditioned policies, one with the GNM dataset alone and one with GNM + in-the-wild videos to evaluate how well MBRA enables us to close the embodiment gap between robot and in-the-wild data.
To evaluate the performance in a variety of situations, we collect the topological memories on four indoor trajectories and four outdoor trajectories and deploy the policies with the ERZ. The distance from the initial node to the goal node is between 10.0 m and 31.0 m. As shown in Table 4, the policy trained with the MBRA-annotated in-the-wild video data has an explicit advantage compared to the policy trained only on the GNM dataset. Although the training dataset does not contain the data from the target robot, ERZ, we achieve a high success rate by training with diverse video data.
5.4 Evaluating MBRA on effectively using crowd-sourced data
To answer Q3, we compare MBRA and multi-step VPT that demonstrated the strongest performance in Table 2. We train several relabelers with different data setups for each method and deploy them as the short-horizon navigation policy in the same eight environments and topological memories as in the previous section to more thoroughly explore the capabilities of each of these relabelers.
Table 5: Comparison of MBRA and multi-step VPT on short-horizon navigation.
Dataset multi-step VPT MBRA
GNM FrodoBots-2k GS SC GS SC
0.500 0.680 0.875 0.960
raw label 0.000 0.308 0.500 0.777
filtered label 0.125 0.377 0.875 0.940
[…]
filtered label 0.375 0.576 1.000 1.000
[…]
Table 6: Evaluation of the goal image-conditioned navigation at 6 countries. The detailed breakdown is shown in appendix.

### v3

@@ IV-B Pre-Processing and Filtering
As shown on the leftmost side of Fig. -1, we use a classical state estimation pipeline to get better coarse robot pose estimates for FrodoBots-2k. We use a smoothing system based on a bidirectional Extended Kalman Filter (EKF) [kalman1960new] to fuse raw actions with wheel speed measurements, GPS location, and compass heading (all of which are noisy) to get a smoothed estimate of the robot’s position. We also filter out data where the robot is paused for a long time to prioritize learning desirable behaviors. The cleaned and filtered data consists of approximately 700 hours of real-world navigation trajectories collected worldwide, which is still an order of magnitude larger than any currently available visual navigation dataset as shown in Table I. While the EKF-based state estimation helps produce a less noisy action estimate [kfsurvey], the signal remains too noisy for direct training.
IV-C Training Details
@@
We describe the training settings for both our short-horizon relabeling model, and long-horizon navigation policy.
Short-horizon relabeling model: Following [hirose2023exaug], to encourage the MBRA model to smoothly connect between $O_{c}$ and $O_{g}$ without collision, we design $s^{s}_{i}$ by three components, $\hat{p}_{i},\hat{c}_{i},\Delta a^{s}_{i}]$, Here $\hat{p}_{i}$ is the $i$-th virtual robot pose, $\hat{c}_{i}$ is the estimated collision state at $i$-th virtual robot pose $\hat{p}_{i}$ (where zero indicates no collision), and $\Delta a^{s}_{i}$ indicates the action difference, $a^{s}_{i+1}-a^{s}_{i}$. Accordingly, we define $s^{ref}$ as $p_{g}$, 0.0, 0.0], where $p_{g}$ is further goal pose.
Since we design $\{a^{s}_{i}\}_{i=0\ldots N-1}$ as the linear and angular velocities, we calculate the unicycle model to integrate the velocity commands for $N$ = 8 steps at 3 Hz and generate the virtual robot poses $\{\hat{p}_{i}\}_{i=0\ldots N-1}$ in the dynamic forward model $f$. In addition, we estimate 3D point clouds from the current image $O_{c}$ via the depth estimation model [hirose2022depth360] and count collided 3D points at $\{\hat{p}_{i}\}_{i=0\ldots N-1}$ as $\{\hat{c}_{i}\}_{i=0\ldots N-1}$ in $f$. By penalizing $\hat{c}_{i}$ to be smaller values, the MBRA model learns to generate collision-free actions. While our objectives are not explicitly designed to learn semantic behaviors such as staying on paths or avoiding pedestrians, the generated trajectories effectively connect $O_{c}$ and $O_{g}$, implicitly capturing such behaviors. Although the action labels are unusable due to heavy noise, the image sequences from teleoperation still reflect the teleoperator’s semantic intent.
@@
Furthermore, since the robot system has $L$ steps system delay [cui2023learning, alnajdi2023machine] when operating remote robot via internet, we design our objective and network architecture to account for system delay to prevent overshooting or oscillating around target trajectories. Inspiring the previous works of model predictive control [kwon2004general, kwon2003simple], we consider the robotic states with the previous action commands $\{a_{i}\}_{i=-L\ldots-1}$ to genrate the actions $\{a_{i}\}_{i=0\ldots N-1}$.
@@
In training, we set the observation and action rate for trajectory sampling at 3 Hz for consistency with the GNM dataset. During training, we randomly select an image frame from the entire dataset as the current observation, and then randomly select a goal frame from up to $N_{g}=$ 20 steps (about 7 seconds) in the future. This short distance to the goal lets us learn precise labels to reannotate the action between $O_{c}$ and $O_{g}$. More details are shown in [hirose2023exaug] and our supplemental code base.
@@
Long-horizon navigation policy: For long-horizon navigation, we use a larger $N_{g}=100$ to sample a goal position up to 33 seconds into the future. We reannotate actions with the short-horizon MBRA model to get high-quality action labels for the FrodoBots-2k dataset. This process yields action labels with a chunk size of $N=8$ steps. We train on the IL objective $J_{il}$ using the same parameters and settings as the short-horizon relabeling model otherwise. Since the action space for long-horizon navigation is the 2D pose, following [shah2023vint], we use the integrated pose commands from the MBRA model as the supervision. In inference, we apply the PD controller to calculate the velocity commands from the generated target pose, similar to [shah2023vint, sridhar2024nomad]. The observation space of $O_{c}$ and $O_{g}$ is the image space for both policies.
Network design: Figure 1 shows the network architecture of both our MBRA model, $\pi^{s}$ and LogoNav policy, $\pi^{l}$. For $\pi^{s}$, we concatenate the current observation $O_{c}$ and the goal observation $O_{g}$ and generate a goal-conditioned embedding with EfficientNet-B0. In addition, we concatenate the image observation history $\{O_{i}\}_{i=-M\ldots 0}$ and generate a history embedding with EfficientNet-B0. We pass in these visual features, the system delay $L$ and the previous action commands $\{a^{s}_{i}\}_{i=-L\ldots-1}$ to a set of Transformer and fully connected MLP layers to produce action commands $\{a^{s}_{i}\}_{i=0\dots N-1}$.
For $\pi^{l}$, we replace the visual encoder for $O_{c}$ and $O_{g}$ with MLP layers for the local goal pose $p_{g}$ on the current robot coordinate in our implementation. In addition, we no longer include system delay length $L$ and previous actions $\{a^{s}_{i}\}_{i=-L\ldots-1}$. Instead of considering the delay during training, we use the $L^{\textrm{th}}$ step of the output during inference, similar to [chi2024universal] and [shah2023vint]. More details are shown in the supplemental code base.
Figure 1: Network architecture. In addition to the visual observations, We feed the delay step and the previous actions to consider the system delay in the MBL objective. For the long-horizon navigation policy, we replace the visual encoder for the current and the goal observation with the MLP layers for the goal pose.
V Evaluation
@@
    Can we use MBRA for action-free in-the-wild data?
@@ V-A Evaluation Setup
We describe both the short-horizon and long-horizon navigation tasks, as well as the robot hardware, on which we evaluate our method, along with their associated baselines.
@@
Short-horizon navigation policy: Short-horizon navigation policy (MBRA model) can navigate the robot toward a goal up to 3 meters away, so we use a topological memory to enable the robot to navigate to further goal positions, similar to other vision-based navigation approaches [savinov2018semi, hirose2019deep, shah2023vint]. To collect this goal loop, we teleoperate the robot and record image observations at a fixed frame rate of 1 Hz. To deploy the policy, we start from the initial observation and continuously estimate the closest node as the current node at each time step, following [shah2023vint, sridhar2024nomad]. We feed the image from the next node as the goal image $O_{g}$ to our policy to compute the next action.
@@
Long-horizon navigation policy: Our LogoNav policy can navigate to goals between 25-100 meters from the initial robot pose in environments unseen during training. We rely on GPS(outdoor) and tracking camera [t265](inside) to get robot positions and specify goals. We evaluate longer trajectories by setting multiple subgoals at intervals of approximately 80 meters apart. At every step, we calculate the relative goal pose $p_{g}$ on the way to the next goal pose. When $|p_{g}|<5.0$ m, we consider the goal reached and update to the next subgoal for a longer trajectory.
Robot hardware: The FrodoBot “Earth Rover Zero” (ERZ), shown in Fig. 2, is a low-cost RC car used both for the FrodoBots-2k dataset collection and our main evaluation. The ERZ includes a host of sensors such as front and back side cameras, GPS, an IMU unit including gyroscope, accelerometer and compass sensors, and wheel velocity sensors in all four wheels. All measurements from the sensors can be accessed through the platform’s API. Linear and angular velocity commands can also be sent to the robot from user teleoperation for data collection (gaming) or from our trained policies for navigation. The rover can turn in place and lasts for five hours on a fully charged battery.
In addition, we conduct additional evaluations with different robot hardware and systems with the wheeled mobile robot, VizBot [niwa2022spatio] and the quadruped robot, Unitree Go1 to analyze the cross-embodiment performance of our policy.
Figure 2: Overview of the robot hardwares and systems. ERZ can be controlled over a internet connection for data collection and for deploying our navigation policy. Vizbot and Go1 with different cameras are controlled from the onboard robot controller with ROS.
V-B Baseline Method
In our evaluation, we use the following two baselines, NoMaD and behavior clolning (BC). For BC, we evaluate various annotation methods as the ground truth action labels to compare with our MBRA model.
NoMaD [sridhar2024nomad]: We deploy the NoMaD policy [sridhar2024nomad] for exploration and generate 30 possible trajectories. Out of these options, we select the best trajectory by measuring the distance between the last predicted position and the goal pose and selecting the minimum one to control the robot.
Behavior Cloning [shah2023vint]: We train a long-horizon navigation policy on reannotated action labels by following several baseline methods instead of using our MBRA. Similar to our methods, we sample 8 steps robot trajectory at 3.0 Hz in the current robot coordinates in all baseline methods. All learning setups except annotation are same as our method.
Raw action label: As the simplest action commands, we annotate the robot trajectory by integrating the teleoperator’s velocity command.
Filtered action label: We give the mentioned EKF for entire FrodoBots-2k dataset to estimate the less-noisy robot pose. We transform them into the local robot coordinate.
Visual SLAM [liu2024citywalker]: Following [liu2024citywalker], we estimate the global trajectories with one of the state-of-the-art viusal SLAM, DPVO [dpvo]. To have better pose estimation, we rectify all images in FrodoBots-2k dataset for DPVO.
VPT [baker2022videopretrainingvptlearning]: We train the inverse dynamics model (IDM) to estimate the relative pose between two consecutive observations such as $p^{i}_{i+1}=f_{idm}(O_{i},O_{i+1})$ by imitating the ground-truth relative pose in the dataset. In training, we sample 9 flames from the current frame to the 8-step future frame and estimate the relative poses between each frame by IDM. Then we integrate the estimated relative poses to have the robot trajectories on the current coordinate.
GCP [shah2023vint]: Following [shah2023vint], we train the policy as GCP to estimate the robot trajectory to link between two frames, $O_{c}$ and $O_{g}$. Since we want to annotate the actions for 8 steps, we select $O_{g}$ as the 8 step future frame from $O_{c}$ in training. The others are same as the original paper [shah2023vint].
For VPT and GCP, we use both the curated GNM dataset and 1 $\%$ FrodoBots-2k dataset to be accurate models. We decide the ratio of the FrodoBots-2k dataset as 1 $\%$ according to the data ablation study in the evaluation section V-E. By mixing small FrodoBots-2k dataset with the clean GNM dataset, VPT and GCP can suppress the negative effect of the noisy FrodoBots-2k dataset and can learn the target robot characteristics. Besides, our MBRA model can use full FrodoBots-2k dataset in training due to the robust learning architecture of the model-based learning.
@@ V-B Baseline Method
TABLE II: Evaluation of LogoNav on long-horizon pose-conditioned navigation tasks. “GS” and “COV” indicate the goal success rate and the coverage rate.
@@
NoMaD [sridhar2024nomad] GNM only - 0.333 0.471
filtered action [kfsurvey] 0.286 0.429
Behavior Cloning raw action 0.286 0.567
filtered action [kfsurvey] 0.286 0.624
visual SLAM [dpvo] 0.286 0.486
VPT [baker2022videopretrainingvptlearning] 0.095 0.314
GCP [shah2023vint] 0.619 0.757
@@
TABLE III: Quantitative analysis with quadruped robot, Go1 and wheeled robot, VizBot for cross-embodiment analysis.
@@
Behavior Cloning GCP 0.300 0.680 0.200 0.630
@@
TABLE IV: Evaluation of MBRA on action-free in-the-wild YouTube videos. “GS” and “SC” indicate the goal success rate and the subgoal coverage rate.
@@
     ✗ 0.500 0.680
     ✓(Visual SLAM [dpvo]) 0.125 0.313
     ✓(MBRA) 0.875 0.909
TABLE V: Evaluation of the goal image-conditioned navigation at six countries.